# Turbulent scaling law in Ogata Kōrin’s Red and White Plum Blossoms

### v1

@@ Abstract
Stylized turbulent swirls depicted in artworks are often analyzed with the modern tools for real turbulent flows such as the power spectrum and the structure function. Motivated by the recent study on The Starry Night of van Gogh (Ma et al., Phys. Fluids, 36 095140, 2024), we here analyze the swirling motif in Ogata Kōrin’s Red and White Plum Blossoms. The results show that Kōrin’s swirls follow closely the Obukhov–Corrsin spectrum $k^{-5/3}$ in the inertial-convective range of the passive scalar advected by the homogeneous and isotropic turbulence. Furthermore the swirl’s 4-th and 6-th structure functions exhibit approximately the same intermittent scaling law of the passive scalar.
@@ I Introduction
Motivated by this recent studyMa et al. (2024), we here analyze stylized turbulent swirls in a Japanese painting in the 18th century utilizing the spectrum and structure function. The painting is Red and White Plum Blossoms by Ogata Kōrin (1658–1716), which is shown in Fig. 1. It is less known than The Great Wave off Kanagawa by Hokusai (1760–1849). Nonetheless, Red and White Plum Blossoms is considered also as one of the most important Japanese artworks of the Edo period (1603–1868). The artists influenced by the style of Kōrin and his predecessors are now called “Rin-pa”Carpenter (2012) coined from his given name. It is known that Hokusai is one of them (see e.g., Refs. Yasui, 1996; Fujisawa, 2022).
Although the title Red and White Plum Blossoms does not mention it, the river is flowing between the red plum tree on the right and the white plum tree on the left. The flow is expressed by boldly stylized swirls that exhibit a irregular hierarchical structure. This intriguing hierarchy leads us to identify its scaling law. In particular, we first show that the luminance spectrum follows closely $k^{-5/3}$, which can be interpreted as the Obukhov–Corrsin spectrumObukhov (1949); Corrsin (1951); Tennekes and Lumley (1972) of the passive scalar advected by the homogeneous and isotropic turbulence. Then we demonstrate that the luminance obeys approximately the same intermittent scaling law of the passive scalar turbulenceFalkovich, Gawȩdzki, and Vergassola (2001); Warhaft (2000); Watanabe and Gotoh (2004); Gotoh and Yeung (2013) by calculating the $4$-th and $6$-th order structure functions. Namely, the power-law exponents of the $p$-th order structure functions of the luminance in the scaling range are different from $p/3$ ($p=4$ and $6$ here), but approximately the same as those of the passive scalar. To the best of our knowledge, such consistency of an artwork with the real turbulent flow including the intermittency is not known.
Why is the stylized flow in Red and White Plum Blossoms created about 300 years ago quantitatively consistent with the detailed laws of the present-day fluid dynamics? Although we do not have an answer to this question yet, we will discuss about this point at the end of this paper together with the insightsDudley, Sarano, and Dias (2013); Ma et al. (2024) obtained from The Great Wave off Kanagawa and The Starry Night.
@@ II Pre-process of the image data
Figure 2: Gray-scaled image of Fig. 1. More specifically, the luminance $\theta(x,y)$ defined in Eq. (2). The annotated circular regions, C1 and C2, are used in the spectral analysis in Sec.III. The radii of C1 and C2 shown here are set to $\sigma$’s listed in Table 1. The square regions, S1 and S2, are used in the structure-function analysis in Sec.IV.
@@ III Spectrum via windowed Fourier transform
Let us calculate the spectrum of $\theta(x,y)$ in the river flowing between the white and red plum trees. Since we focus only on the river part of the painting, we adopt the windowed Fourier transform, also known as the Gabor transform. We consider two regions for calculating the spectrum. One of them is taken in the left panel as depicted C1 in Fig. 2 and the other one is taken in the right panel as depicted C2 in the same figure. The windowing is done with a two-dimensional Gaussian function,
@@
The parameter values for C1 and C2 are listed in Table 1. The windowed luminance field $\theta_{w}(x,y)$ for the two regions are shown in Fig. 3. Indeed they include only the river regions.
Figure 3: Gaussian-windowed gray scale images for calculating the spectrum. The top and bottom panels correspond respectively to C1 and C2 in Fig. 2. The black color represents zero luminance after multiplied by the Gaussian function. The non-zero luminance (gray or white) in this figure is enhanced than that of Fig. 2 to highlight the effect of the windowing.
@@
The resultant spectra for C1 and C2 are plotted in Fig. 4. We observe the power-law behavior in the intermediate wavenumber range for both cases. This power law is consistent with the Obukhov–Corrsin (OC) spectrum, $C\chi\epsilon^{-1/3}k^{-5/3}$, in the inertial-convective range for the passive scalar advected by HITObukhov (1949); Corrsin (1951); Tennekes and Lumley (1972). Here $C$, $\chi$, and $\epsilon$ are the non-dimensional Obukhov–Corrsin constant, the scalar dissipation rate, and the energy dissipation rate, respectively.
This $k^{-5/3}$ like power-law part of the luminance spectrum is followed by the steeply decreasing part in the larger wavenumbers. This is in contrast to the spectrum found in The Starry Night where the $k^{-5/3}$ part is followed by the shallower Batchelor spectrum $k^{-1}$ of the dissipative-convective range Ma et al. (2024). Our decreasing part for $k\geq 1$ [(pixel)-1] in Fig. 4 may correspond to the dissipation-range behavior (non power-law behavior) of the passive scalar spectrum. Or it may correspond to the Batchelor–Howells–Townsend spectrum, $\tilde{C}\chi\kappa^{-3}\epsilon^{2/3}k^{-17/3}$, of the passive scalar at low Schmidt numbers in the inertial-diffusive range Batchelor, Howells, and Townsend (1959), as indicated in Fig. 4 by the dashed line. Here $\tilde{C}$ and $\kappa$ are the non-dimensional constant and the diffusivity of the passive scalar, respectively. The two possibilities suggest that the luminance is analogous to unit or low Schmidt number cases. However, the range $k\geq 1~{}\text{[(pixel)]}^{-1}$ is too narrow to conclude which one is more plausible. Furthermore, the smallest scales of the image can be more affected by the image compression or other procedures of the image processing. In this paper, we focus on the consistency with the OC spectrum.
In Fig. 4 we do not show the luminance spectrum in the small wavenumber range. That part of the spectrum is the Gaussian function, namely $E_{\theta}(k)\propto\exp(-\sigma^{2}k^{2})$, due to the windowing. The flat part of the luminance spectrum shown in Fig. 4 ends roughly at $k=2\pi/\sigma$ and the appearance of the power-law part in $k>2\pi/\sigma$ indicates that the $k^{-5/3}$-like power law reflects the nature of Kōrin’s flowing water motif. Given this consistency with the OC scaling law, we now address whether the intermittency corrections present, although the scaling range of the $k^{-5/3}$-like part of the spectrum in Fig. 4 is not more than a decade. We calculate the structure functions of the luminance in the next section.
@@
Figure 4: Spectra of the luminance calculated in the two circular regions, C1 and C2, with the windowed Fourier transform. The solid line corresponds to the Obukuhov–Corrsin spectrum, $k^{-5/3}$, in the inertial-convective range for the passive scalar advected by the homogeneous and isotropic turbulence. The dashed line corresponds to the Batchelor–Howells–Townsend spectrum, $k^{-17/3}$, in the inertial-diffusive range for the passive scalar turbulence. Here $1$ [pixel] corresponds to $0.089$ [cm].
@@ IV Structure function
Figure 5: Even-order structure functions, $S_{2}(r)$ (a), $S_{4}(r)$ (b), and $S_{6}(r)$ (c), of the luminance calculated in the two square regions S1 and S2. The solid lines denote the power-law exponents of the structure functions of the passive scalar. The exponent $2/3$ for the second-order one corresponds to the OC spectrum $k^{-5/3}$, for which the intermittency corrections are ignored. The values of the other exponents ($1.04$ for $p=4$ and $1.29$ for $p=6$) are taken from the numerical study of the passive scalar in the homogeneous and isotropic turbulence Watanabe and Gotoh (2004), so that the intermittency corrections are present. As used in Fig. 4, one pixel corresponds to $0.089$ cm.
@@
The structure functions for $p=2,4$ and $6$ calculated in this manner are shown in Fig. 5. We here focus on those even-order functions since they converge more quickly than odd-order ones. For the range of $r$ where we expect a power-law scaling, let us go back to Fig. 4 in which the luminance spectrum is close to the OC scaling roughly in $0.2\leq k\leq 1$ (with the unit [$(\text{pixel})^{-1}$]). This wavenumber range corresponds to $6\leq r\leq 30$ (with the unit [pixel]).
@@
However, as shown in Fig. 5(b) and (c), the luminance structure functions for $p=2$ and $4$ differ from $r^{4/3}$ and $r^{2}$, respectively. Instead they are closer to $r^{1.04}$ and $r^{1.29}$. Those values of the exponents Watanabe and Gotoh (2004) are so-called intermittent scaling exponents observed for the structure functions of the passive scalar advected in HIT. By intermittentFrisch (1995) we mean that the exponent of $S_{p}(r)$ is different from $p/3$ that is obtained by the dimensional argumentObukhov (1949); Corrsin (1951); Tennekes and Lumley (1972) leading to the OC spectrum $k^{-5/3}$. It is known that, for the passive scalar turbulence, the power-law exponent for $p=2$ differs from $2/3$ only slightly. However, for $p>2$, the difference becomes larger, a phenomenon known as the intermittencyFrisch (1995); Falkovich, Gawȩdzki, and Vergassola (2001); Warhaft (2000); Gotoh and Yeung (2013). Those differences, called intermittency corrections, are highlighted in the flatness $S_{4}(r)/S_{2}^{2}(r)$ and the hyperflatness $S_{6}(r)/S_{2}^{3}(r)$ of the luminance increment, as shown in Fig. 6. Variations are present depending on the regions, S1 and S2. But, overall, their $r$-dependence in the scaling range indicates that the intermittency of the luminance is approximately the same as the passive scalar advected by HIT.
@@
The intermittency of the structure functions can be a manifestation of the multifractalityFrisch (1995) of the luminance. It implies that the scale invariance of the luminance is broken. This can be observed in the probability distribution function (PDF) of the luminance increment for different $r$’s. We show the PDFs in the S1 and S2 regions in Fig. 7 for the normalized increment $\widetilde{\delta_{r}\theta}=[\delta_{r}\theta-\langle\delta_{r}\theta\rangle]/[\langle(\delta_{r}\theta)^{2}\rangle-\langle\delta_{r}\theta\rangle^{2}]^{1/2}$. Obviously the PDFs are not Gaussian, having fat tails. More importantly, the normalized PDFs for several $r$’s in the scaling range are not collapsed to one master curve, although the differences between the PDFs are small. Thus the scale invariance being broken. This non-collapsed behavior of the non-Gaussian PDF is similar to that of the passive scalar in HITWarhaft (2000); Watanabe and Gotoh (2004); Gotoh and Yeung (2013).
In summary, we find that Kōrin’s flowing water motif respects approximately the physical law of the passive scalar advected by HIT. By the law, we mean not only the OC scaling of the second order quantities, but also the intermittent exponents of the fourth and sixth order structure functions. Those intermittent exponents of the passive scalar turbulence are known to be universal in the sense that they do not depend on details of the system, such as how to inject the scalar quantity into the turbulent flow Falkovich, Gawȩdzki, and Vergassola (2001); Gotoh and Yeung (2013). Such agreement for this highly stylized depiction of flow is likely not trivial at all and thus comes as a delightful surprise to us.
Table 2: Coordinates of the square regions, S1 and S2, for the structure functions. The centers are the same as those of C1 and C2 used in the windowed Fourier transform.
@@
Refer to captionRefer to caption
Figure 7: Probability distribution functions of the luminance increments in the region S1 (a) and the region S2 (b) for several separation $r$’s in the scaling range of the structure functions. The values of $r$’s are given in pixels. The increment denoted by $\widetilde{\delta\theta}$ here is the normalized increment such that its mean and standard deviation are zero and unity, respectively.

### v2

@@ Abstract
Stylized turbulent swirls depicted in artworks are often analyzed with the modern tools for real turbulent flows such as the power spectrum and the structure function. Motivated by the recent study on The Starry Night of van Gogh (Ma et al., Phys. Fluids, 36 095140, 2024), we here analyze Ogata Kōrin’s Red and White Plum Blossoms, in particular its swirling pattern and the bark of the plum-tree trunk. The results show that they follow closely the Obukhov–Corrsin spectrum $k^{-5/3}$ in the inertial-convective range of the passive scalar advected by the homogeneous and isotropic turbulence. Furthermore their 4th- and 6th-order structure functions exhibit approximately the same intermittent scaling law of the passive scalar. We discuss several possible explanations of this consistency.
@@ I Introduction
Motivated by this recent studyMa et al. (2024), we here analyze stylized turbulent swirls and another part in a Japanese painting in the 18th century by utilizing the spectrum and structure function. The painting is Red and White Plum Blossoms by Ogata Kōrin (1658–1716), which is shown in Fig. 1. It is less known than The Great Wave off Kanagawa by Hokusai (1760–1849). Nonetheless, Red and White Plum Blossoms is considered also as one of the most important Japanese artworks of the Edo period (1603–1868). The artists influenced by the style of Kōrin and his predecessors are now called “Rin-pa”Carpenter (2012) coined from his given name. It is known that Hokusai is one of them (see e.g., Refs. Yasui, 1996; Fujisawa, 2022).
Although the title Red and White Plum Blossoms does not mention it, the river is flowing between the red plum tree on the right and the white plum tree on the left. The flow is expressed by boldly stylized swirls that exhibit a irregular hierarchical structure. This intriguing hierarchy leads us to identify its scaling law. Furthermore we analyze also the bark of the trunk of the white plum tree. In particular, we first show that the luminance spectrum follows closely $k^{-5/3}$, which can be interpreted as the Obukhov–Corrsin spectrumObukhov (1949); Corrsin (1951); Tennekes and Lumley (1972) of the passive scalar advected by the homogeneous and isotropic turbulence. Then we demonstrate that the luminance obeys approximately the same intermittent scaling law of the passive scalar turbulenceFalkovich, Gawȩdzki, and Vergassola (2001); Warhaft (2000); Watanabe and Gotoh (2004); Gotoh and Yeung (2013) by calculating the $4$th- and $6$th-order structure functions. Namely, the power-law exponents of the $p$-th order structure functions of the luminance in the scaling range are different from $p/3$ ($p=4$ and $6$ here), but approximately the same as those of the passive scalar. To the best of our knowledge, such consistency of an artwork with the real turbulent flow including the intermittency is not known. For the stylized swirls, it may appear conceivable, but the same consistency holds also for the the depicted bark pattern on the plum tree.
Why are the flow pattern and the bark pattern on the plum tree in Red and White Plum Blossoms created about 300 years ago quantitatively consistent with the detailed laws of the present-day fluid dynamics? Is it just a coincidence? Although we do not have an answer to this question yet, we will discuss several possible explanations at the end of this paper together with the insightsDudley, Sarano, and Dias (2013); Ma et al. (2024) obtained from The Great Wave off Kanagawa and The Starry Night.
@@ II Pre-process of the image data
Figure 2: Gray-scaled image of Fig. 1. More specifically, the luminance $\theta(x,y)$ defined in Eq. (2). The annotated circular regions, C1, C2, and C3, are used in the spectral analysis in Sec.III. The radii of C1, C2, and C3 shown here are set to $\sigma$’s listed in Table 1. The square regions, S1, S2, and S3, are used in the structure-function analysis in Sec.IV.
@@ III Spectrum via windowed Fourier transform
Let us calculate the spectrum of $\theta(x,y)$ in several parts of the painting. We consider the parts including bulk of the river flowing between the white and red plum trees and the part other than the river, namely the bark of the white plum-tree trunk. Since we focus on local parts of the painting, we adopt the windowed Fourier transform, also known as the Gabor transform. We consider three regions for calculating the spectrum. One of them is taken in the left panel of the screen as depicted C1 in Fig. 2 and the second one is taken in the right panel as depicted C2 in the same figure. The third one is taken in the left panel as marked C3 in in the same figure. The regions C1 and C2 are in the river part and C3 is on the trunk of the white plum. The windowing is done with a two-dimensional Gaussian function,
@@
The parameter values for C1, C2 and C3 are listed in Table 1. The windowed luminance field $\theta_{w}(x,y)$ for the three regions are shown in Fig. 3. Indeed they highlight the local regions.
Figure 3: Gaussian-windowed gray scale images for calculating the spectrum. The top, middle and bottom panels correspond respectively to C1, C2 and C3 in Fig. 2. The black color represents zero luminance after multiplied by the Gaussian function. The non-zero luminance (gray or white) in this figure is enhanced than that of Fig. 2 to highlight the effect of the windowing.
@@
The resultant spectra for C1, C2 and C3 are plotted in Fig. 4. We observe the power-law behavior in the intermediate wavenumber range for all three cases. This power law is consistent with the Obukhov–Corrsin (OC) spectrum, $C\chi\epsilon^{-1/3}k^{-5/3}$, in the inertial-convective range for the passive scalar advected by HITObukhov (1949); Corrsin (1951); Tennekes and Lumley (1972). Here $C$, $\chi$, and $\epsilon$ are the non-dimensional Obukhov–Corrsin constant, the scalar dissipation rate, and the energy dissipation rate, respectively.
This $k^{-5/3}$ like power-law part of the luminance spectrum is followed by the steeply decreasing part in the larger wavenumbers. This is in contrast to the spectrum found in The Starry Night where the $k^{-5/3}$ part is followed by the shallower Batchelor spectrum $k^{-1}$ of the dissipative-convective range Ma et al. (2024). The decreasing part for $k\geq 1$ [(pixel)-1] in Fig. 4 may correspond to the dissipation-range behavior (non power-law behavior) of the passive scalar spectrum. Or it may correspond to the Batchelor–Howells–Townsend spectrum, $\tilde{C}\chi\kappa^{-3}\epsilon^{2/3}k^{-17/3}$, of the passive scalar at low Schmidt numbers in the inertial-diffusive range Batchelor, Howells, and Townsend (1959), as indicated in Fig. 4 by the dashed line. Here $\tilde{C}$ and $\kappa$ are the non-dimensional constant and the diffusivity of the passive scalar, respectively. The two possibilities suggest that the luminance is analogous to unit or low Schmidt number cases. However, the range $k\geq 1~{}\text{[(pixel)]}^{-1}$ is too narrow to conclude which one is more plausible. Furthermore, the smallest scales of the image can be more affected by the image compression or other procedures of the image processing. In this paper, we focus on the consistency with the OC spectrum.
In Fig. 4 we do not show the luminance spectrum in the small wavenumber range. That part of the spectrum is the Gaussian function, namely $E_{\theta}(k)\propto\exp(-\sigma^{2}k^{2})$, due to the windowing. The flat part of the luminance spectrum shown in Fig. 4 ends roughly at $k=2\pi/\sigma$ and the appearance of the power-law part in $k>2\pi/\sigma$ indicates that the $k^{-5/3}$-like power law reflects the nature of Kōrin’s flowing water motif and pattern of the bark. Now we observed that the spectra of the painting shown in Fig. 4 are close to $k^{-5/3}$. However we cannot rule out that this consistency with the OC scaling law is just a coincidence, since the pattern of the bark is not related to fluid motion. In the next section, we seek a different kind of consistency by calculating the even-order structure functions of the luminance up to 6th order.
@@
C3 $(280,1275)$ $35$
Figure 4: Spectra of the luminance calculated in the three circular regions, C1, C2, and C3, with the windowed Fourier transform. The solid line corresponds to the Obukuhov–Corrsin spectrum, $k^{-5/3}$, in the inertial-convective range for the passive scalar advected by the homogeneous and isotropic turbulence. The dashed line corresponds to the Batchelor–Howells–Townsend spectrum, $k^{-17/3}$, in the inertial-diffusive range for the passive scalar turbulence. Here $1$ [pixel] corresponds to $0.089$ [cm].
@@ IV Structure function
Figure 5: Even-order structure functions, $S_{2}(r)$ (a), $S_{4}(r)$ (b), and $S_{6}(r)$ (c), of the luminance calculated in the three square regions S1, S2 and S3. The solid lines denote the power-law exponents of the structure functions of the passive scalar. The exponent $2/3$ for the second-order one corresponds to the OC spectrum $k^{-5/3}$, for which the intermittency corrections are ignored. The values of the other exponents ($1.04$ for $p=4$ and $1.29$ for $p=6$) are taken from the numerical study of the passive scalar in the homogeneous and isotropic turbulence Watanabe and Gotoh (2004), so that the intermittency corrections are present. As used in Fig. 4, one pixel corresponds to $0.089$ cm.
@@
The structure functions for $p=2,4$ and $6$ calculated in this manner are shown in Fig. 5. We here focus on those even-order functions since they converge more quickly than odd-order ones (the third-order function with and without the absolute value is discussed in the Appendix A). For the range of $r$ where we expect a power-law scaling, let us go back to Fig. 4 in which the luminance spectrum is close to the OC scaling roughly in $0.2\leq k\leq 1$ (with the unit [$(\text{pixel})^{-1}$]). This wavenumber range corresponds to $6\leq r\leq 30$ (with the unit [pixel]), which are between the two vertical dotted lines in Fig. 5.
@@
However, as shown in Fig. 5(b) and (c), the luminance structure functions for $p=2$ and $4$ differ from $r^{4/3}$ and $r^{2}$, respectively. Instead they are closer to $r^{1.04}$ and $r^{1.29}$. Those values of the exponents Watanabe and Gotoh (2004) are so-called intermittent scaling exponents observed for the structure functions of the passive scalar advected in HIT. By intermittentFrisch (1995) we mean that the exponent of $S_{p}(r)$ is different from $p/3$ that is obtained by the dimensional argumentObukhov (1949); Corrsin (1951); Tennekes and Lumley (1972) leading to the OC spectrum $k^{-5/3}$. It is known that, for the passive scalar turbulence, the power-law exponent for $p=2$ differs from $2/3$ only slightly. However, for $p>2$, the difference becomes larger, a phenomenon known as the intermittencyFrisch (1995); Falkovich, Gawȩdzki, and Vergassola (2001); Warhaft (2000); Gotoh and Yeung (2013). Those differences, called intermittency corrections, are highlighted in the flatness $S_{4}(r)/S_{2}^{2}(r)$ and the hyperflatness $S_{6}(r)/S_{2}^{3}(r)$ of the luminance increment, as shown in Fig. 6. Variations are present depending on the regions, S1, S2 and S3. But, overall, their $r$-dependence in the scaling range indicates that the intermittency of the luminance is approximately the same as the passive scalar advected by HIT.
@@
The intermittency of the structure functions can be a manifestation of the multifractalityFrisch (1995) of the luminance. It implies that the scale invariance of the luminance is broken. This can be observed in the probability distribution function (PDF) of the luminance increment for different $r$’s. We show the PDFs in the S1, S2 and S3 regions in Fig. 7 for the normalized increment $\widetilde{\delta_{r}\theta}=[\delta_{r}\theta-\langle\delta_{r}\theta\rangle]/[\langle(\delta_{r}\theta)^{2}\rangle-\langle\delta_{r}\theta\rangle^{2}]^{1/2}$. Obviously the PDFs are not Gaussian, having fat tails. More importantly, the normalized PDFs for several $r$’s in the scaling range are not collapsed to one master curve, although the differences between the PDFs are small. Thus the scale invariance is broken. This non-collapsed behavior of the non-Gaussian PDF is similar to that of the passive scalar in HITWarhaft (2000); Watanabe and Gotoh (2004); Gotoh and Yeung (2013). We notice that the tail of the PDF for S3 (the bark) in Fig. 7(c) is less pronounced than those for S1 and S2 (the swirls). This tail difference does not result in sizable differences in the structure functions up to the 6th order shown in Fig. 5.
In summary, we find that Kōrin’s flowing water motif and the bark of the plum-tree trunk respect approximately the physical law of the passive scalar advected by HIT. By the law, we mean not only the OC scaling of the second order quantities, but also the intermittent exponents of the 4th- and 6th-order structure functions. Those intermittent exponents of the passive scalar turbulence are known to be universal in the sense that they do not depend too much on details of the system, such as how to inject the scalar quantity into the turbulent flow Falkovich, Gawȩdzki, and Vergassola (2001); Gotoh and Yeung (2013). Such agreement for this highly stylized depiction of flow (S1 and S2) is likely not trivial at all and thus comes as a delightful surprise to us. The agreement for the bark (S3) appears puzzling since it is not a depiction of a flow. In the next section we will argue as one possibility that it has something to do with the painting technique used for the bark.
Table 2: Coordinates of the square regions, S1, S2 and S3, for the structure functions. The centers are the same as those of C1, C2 and C3 used in the windowed Fourier transform.
@@
Refer to captionRefer to captionRefer to caption
Figure 7: Probability distribution functions of the luminance increments in the region S1 (a), S2 (b), and S3 (c) for several separation $r$’s in the scaling range of the structure functions. The values of $r$’s are given in pixels. The increment denoted by $\widetilde{\delta\theta}$ here is the normalized increment such that its mean and standard deviation are zero and unity, respectively.